OpenStudy (anonymous):

What's the area of an equilateral triangle whose perimeter is 24 centimeters

OpenStudy (campbell_st):

so if the perimeter is 24, and an equilateral triangle has all sides equal. then each side is 24/3 = 8 centimetres. so now you need to find perpendicular height using pythagoras' theorem  find h then you can find the area. hope it helps

OpenStudy (anonymous):

Still confused :(

OpenStudy (campbell_st):

ok... do you know about pythagoras theorem...?

OpenStudy (campbell_st):

otherwise.... wait and someone always gives the answer

OpenStudy (koikkara):

@emmadeezy Hello, u have to answer him because OS appreciates only "how to solve it" not "direct Answers"! So, what u want me to do, explain ??

OpenStudy (anonymous):

@emmadeezy, when you're back online, tag me, and I'll help you.

OpenStudy (anonymous):

Ok, building off what campbell_st said, each side is 8 cm. Using the Pythagorean Theorem, the length of the squared hypotenuse-the squared base=the square root of the height. \[8^{2}=64\]\[4^{2}=16\]\[64-16=48\]\[\sqrt{48}=6.92\]So, the height is 6.92 cm. The area of a triangle is; A=.5*h*b\[.5\times6.92\times8=27.71\]So, the area of the equilateral triangle is 27.71 cm. ^2

OpenStudy (anonymous):

Is that all you needed? @emmadeezy